### v7

@@
To implement LV, we need to interpolate and extrapolate option prices. This approach is significantly faster and more accurate than any parameterized model. The implementation is demonstrated specifically for interest rate swaptions. A key component of this method is the Small Volatility Approximation within the HJM interest rate model, which is used to calculate forward volatility sensitivities. These sensitivities are deterministic and fast, with excellent calibration accuracy.
@@ 1 Introduction
The simplest way to do it is to assume that smile has a quadratic form for all points on the grid.
As we can see from Fig.4 this assumption does not work very well in case of short-term smiles.
However in case of longer smiles this assumption works good (see Fig.5).
@@
Using this procedure, we can compute all forward rates and, consequently, the corresponding swaptions. The results for ATM swaptions are shown in Figures 6, and the calibration quality appears to be good.
The volatility smiles also look satisfactory, as seen in Figures 7β10. However, the 30-year tenor with 30 years to expiration does not exhibit sufficient quality.
@@
Here, we present the implementation of the HJM Local Volatility Model. The results show excellent agreement with input swaption prices across a wide range of expiration dates and tenors. This model is suitable for the pricing of all path-independent interest rate derivatives.

### v8

@@
To implement LV, we need to interpolate and extrapolate option prices. This approach is significantly faster and more accurate than any parameterized model. The implementation is demonstrated specifically for interest rate swaptions and caplets.
A key component of this method is the Small Volatility Approximation within the HJM interest rate model, which is used to calculate sensitivity of forward bond volatility. These calculations are deterministic and fast, with excellent calibration accuracy.
@@
The simplest way to do it is to assume that implied variance has a quadratic form for all points on the grid:
where is implied total variance; .
As we can see from Figs.4 this assumption does not work very well in case of short-term smiles but
in case of longer smiles this assumption works good.
@@
Using this procedure, we can compute all forward rates and, consequently, the corresponding swaptions. The results for ATM swaptions are shown in Figs.5, and the calibration quality appears to be good.
The volatility smiles also look satisfactory, as seen in Figs.6-9. However, the 30-year tenor with 30 years to expiration does not exhibit sufficient quality.
7 Caplets and Floorlets.
Described procedure can also been used to be calibrated to all caplets and floorlets in addition to swaptions.
Distribution of present values of Caplet is
where is caplet rate; is a standard normal distributed stochastic variable;
and
where caplet market implied volatility is:
We can calculate caplet prices without caplet market and with caplet input.
If we do not use caplets in calibration process then ATM caplet prices are in relative good agreement with market data, but they have significantly lower quality compare to swaption prices Figs.10.
To be able generate LV scenarios calibrated to caplets we need to build model for caplet sensitivities. Here we use the same quadratic model as in case of swaptions.
As we can see from Figs.11 interpolation of 6 months smile does not look good, but 1 year smile is more acceptable.
Using these interpolation smiles we can calculate forward bond sensitivities the same way as it was done in case of swaptions.
ATM caplets looks better than without caplet calibration Fig.12. Short term error can be explained by bad smile interpolation.
The same type of quality we observed for shifted strikes Figs.13. Quality of shifted swaption calibration is the same as in case of
swaption calibration only.
Modeled caplet smiles look also good (see Fig.15).
@@
Here, we present the implementation of the HJM Local Volatility Model.
The results show excellent agreement with input swaption prices across a wide range of expiration dates and tenors.
Caplet calibration is also in a good agreement with market data.
This model is suitable for the pricing of all path-independent interest rate derivatives.